TGBT14A_912_91 - Continuing tests


  • Continue tests from yesterday


  • 8:00am - switch VEGAS versions, etc
  • Configure for all eight banks of VEGAS
  • bank_gain_mode1_xband (all 8 banks)
  • archivist
  • balance
  • scans 1-3: track - results look "normally bad"
  • Galen terminates all of the "0" inputs (A0,B0...H0) at inputs to VEGAS IF modules
  • No other rebalancing or anything.
  • scans 4-6: track.
  • set up with ADC Bank A = A0,B0, B=A1,B1, C=C0,D0,D=C1,D1... same for E through H.
  • Doors now closed
  • scans 7-9. Track
  • Open doors
  • scan 10-12. Track - looks similar to 7-9
  • Terminate all "1" inputs (A1,B1...H1) at inputs to VEGAS IF modules
  • Scans 13-15. Track
  • Put one signal into each of eight IF modules, with other input terminated.
  • ADC banks A,C,E,H, A = A0,A1, C=C0,C1, etc...
  • Scan 16-18
  • As above, but flip optical receiver paths in cabling.
  • A swapped with E, C swapped with G.
  • Scans 19-21
  • Tsys switches as expected
  • CM1 - switch OR1 to OR2, same in CM2,3,4. In IF Rack, switch S9. need to set filter and attenuator for this OR path, but cannot exceed 3V
  • Need to set CM1 and CM3 attenuators to remove attenuation
  • This switches a single optical receiver path, but uses the same converter modules (as confirmed by Galen!)
  • Scans 22-24 - all lines up.
  • Switch back to 19-21 configuration. Did NOT need to change converter rack attenuation, implying OR attenuator was previously wrong.
  • Scan 25-27 - all line up

  • Conclusion - Tsys offsets above seem to have been caused by an over-driven optical receiver, which has been there since the very first balance. So, we believe that has gone away.

    • ADCs A,C,E,G have their correct configurations. (A0 is connected as normal. A1 goes to B0, then back to A1 ADC).

  • cabling file checks: create a bunch of "_fixed" scripts, which configure for 4 banks. Make sure everything works.
  • scan 27 - bank_gain_mode1_xband_fixed, banks hard coded. Looks fine. Take out hard coding, and everything works as expected.
  • 28: mode 4. Tsys = 28.81,28.91,26.36,26.37; bandpass shapes different. Dave's routines won't work on these - not sure why not.
  • 29: mode 7. Fails to do internal requantization balancing - perhaps due to inconsistent BOF file / Manager?
  • Since we don't have all new BOFs, need to switch back to "production" manager anyway.
  • 30: mode 4 (old BOF). - ok
  • 31: mode 1 - ok.
  • 32: mode 7 - configures and balances ok, but aborts with an "invalid data" error. Cycle VEGAS Off and On. Seems to fix the problem.
  • 33: mode 7 - ok
  • 34: mode 19 - ok. Consistent Tsys values to 0.05K /0.01K within same pols, 2.0K between Pols (7%)

  • We believe we have found the "cross-talk" problem.
  • Astronomical tests to do CICADAGreenBankSpectrometerTsysVariationsResolved:
    • Make sure the cabling file modification allows us to configure correctly.
    • Repeat the "good" configuration tests on every different BOF combination: Mode 1,2,4,10, the Balser mode (can't remember it now, 20 or 21)
    • Repeat the "good" configuration on a strong flux calibrator in the above four modes.
    • Repeat the "good" configuration with Dana's RRL setup.

-- RichardPrestage - 2015-04-15
Topic revision: r6 - 2015-04-20, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback